# The Fourier spectral approach to the spatial discretization of quasilinear hyperbolic systems

### v1

@@ Appendix A Technical tools
The following result is shown on $\mathbb{R}^{d}$ in [8, Lemma 4.5]. It is straightforward to adapt the proof to the periodic setting.
@@
Notice that if $G$ is such that $\big{|}G\big{|}_{L^{\infty}},\big{|}\left|\,\cdot\,\right|\nabla G\big{|}_{L^{\infty}}\leq C_{G}$, then $G_{N}\coloneq G(\cdot/N)$ satisfies
$\big{|}G_{N}\big{|}_{L^{\infty}},\big{|}\left|\,\cdot\,\right|\nabla G_{N}\big{|}_{L^{\infty}}\leq C_{G}$
uniformly with respect to $N$.
@@ Acknowledgements
VD thanks the program Centre Henri Lebesque ANR-11-LABX-0020-0 for fostering an attractive mathematical environement. JUM acknowledges the support of the project IMod (Grant No. 325114) from the Research Council of Norway.

### v2

@@ Appendix A Technical tools
The following result is shown on $\mathbb{R}^{d}$ in [8, Lemma 4.5]. The proof straightforwardly adapts to the periodic setting.
@@
Notice that for smooth symbols considered in this work, namely $S_{N}(\cdot)=S(\cdot/N)$ where $S$ is even with
$\begin{cases}S(\bm{k})=1&\text{if $\max_{j=1,\ldots,d}|k_{j}|\leq 1/2$,}\\ S(\bm{k})=0&\text{if $\min_{j=1,\ldots,d}|k_{j}|\geq 1$,}\\ S(\bm{k})\in{[}0,1{]}&\text{otherwise,}\end{cases}$
and $S^{1/2}$ is Lipschitz-continuous, $S_{N}^{1/2}$ satisfies the hypotheses of Proposition A.6 uniformly with respect to $N$. Indeed, by Rademacher’s theorem we have that $S_{N}^{1/2}$ is differentiable almost everywhere and its derivative is essentially bounded, and since $S$ has compact support, $\left|\,\cdot\,\right|\nabla S_{N}^{1/2}\in L^{\infty}$. Moreover, we have
$\big|S_{N}^{1/2}\big|_{L^{\infty}}+\big|\left|\,\cdot\,\right|\nabla S_{N}^{1/2}\big|_{L^{\infty}}=\big|S^{1/2}\big|_{L^{\infty}}+\big|\left|\,\cdot\,\right|\nabla S^{1/2}\big|_{L^{\infty}}.$
@@ Acknowledgements
VD thanks Centre Henri Lebesque ANR-11-LABX-0020-0 for fostering an attractive mathematical environement. JUM acknowledges the support of the project IMod (Grant No. 325114) from the Research Council of Norway.